# VideoLLaMA 3: Frontier Multimodal Foundation Models for Image and Video Understanding

### v2

@@ Abstract
In this paper, we propose VideoLLaMA3, a more advanced multimodal foundation model for image and video understanding. The core design philosophy of VideoLLaMA3 is vision-centric. The meaning of “vision-centric” is two-fold: the vision-centric training paradigm and vision-centric framework design. The key insight of our vision-centric training paradigm is that high-quality image-text data is crucial for both image and video understanding. Instead of preparing massive video-text datasets, we focus on constructing large-scale and high-quality image-text datasets. VideoLLaMA3 has four training stages: 1) Vision Encoder Adaptation, which enables vision encoder to accept images of variable resolutions as input; 2) Vision-Language Alignment, which jointly tunes the vision encoder, projector, and LLM with large-scale image-text data covering multiple types (including scene images, documents, charts) as well as text-only data. 3) Multi-task Fine-tuning, which incorporates image-text SFT data for downstream tasks and video-text data to establish a foundation for video understanding. 4) Video-centric Fine-tuning, which further improves the model’s capability in video understanding. As for the framework design, to better capture fine-grained details in images, the pretrained vision encoder is adapted to encode images of varying sizes into vision tokens with corresponding numbers, rather than a fixed number of tokens. For video inputs, we reduce the number of vision tokens according to their similarity so that the representation of videos will be more precise and compact. Benefit from vision-centric designs, VideoLLaMA3 achieves compelling performances in both image and video understanding benchmarks.
@@ 1 Introduction
Figure 1: Performance Comparison of VideoLLaMA3 with the previous advanced image/video MLLM on various representative benchmarks. As shown in the figure, VideoLLaMA3 has achieved very competitive results on various benchmarks. Specifically, VideoLLaMA3 not only demonstrates strong video understanding capabilities (VideoMME, PerceptionTest, MLVU) but also maintains excellent document comprehension abilities (DocVQA) and multimodal mathematical reasoning skills (MathVista). Note that LLaVA-OneVision is only used for evaluating image benchmarks, while LLaVA-Video is only used for evaluating video benchmarks.
@@
Existing MLLMs  lin2023video ; ataallah2024minigpt4 ; zhang2024longva ; zohar2024apollo ; song2024moviechat+ ; fei2024video ; liu2024kangaroo ; liu2024oryx ; wang2024internvideo2 ; jung2024pegasus ; Maaz2024VideoGPT+ ; cheng2024videollama ; sun2024video ; zhang2024video ; Maaz2023VideoChatGPT ; li2023mvbench ; chen2024sharegpt4video ; li2024llavaonevision ; mplug3 ; qwen2-vl ; chen2023internvl ; internvl2.5 ; longllava ; internlmxcomposer25 have made significant progress in multimodal understanding. Image-centric MLLMs qwen2-vl ; li2024llavaonevision ; internvl2.5 ; chen2023internvl ; molmo2024 ; liu2024nvila ; smolvlm2023 , leveraging high-quality image-text datasets li2024llava ; li2024llavaonevision ; tong2024cambrian ; mathew2021docvqadatasetvqadocument ; laurençon2024building ; masry2022chartqa ; huang-etal-2024-chart that are easier to collect and curate, have demonstrated strong performance in image understanding, such as visual question answering, OCR, and document understanding. Beyond static content like images, video-centric MLLMs qwen2-vl ; internvl2.5 ; zhang2024video ; chen2024sharegpt4video ; cheng2024videollama ; damonlpsg2023videollama must tackle the added complexity of modeling the temporal dimension of videos, requiring them to handle dynamic content and capture dependencies across frames. This temporal complexity, combined with the need for large-scale video-text datasets that are often of lower quality and harder to annotate, makes video MLLMs more challenging. These challenges underscore the advantages of leveraging image understanding as a foundation for video understanding. By extending the robust visual capabilities of image MLLMs, video models can focus on and better address the unique challenges of temporal and dynamic content modeling.
@@
Inherit from VideoLLaMA DBLP:conf/emnlp/ZhangLB23 and VideoLLaMA2 damonlpsg2024videollama2 , VideoLLaMA3, a more advanced multimodal foundation model, is proposed for both image and video understanding. We design VideoLLaMA3 in a vision-centric way. Specifically, we propose a vision-centric training paradigm and vision-centric framework designs. For the training paradigm, considering the intrinsic relationship between image and video modalities - where videos are essentially sequences of temporally correlated images, we prioritize the improvement of image understanding, which in turn enhances the performance of video understanding. Moreover, compared to video-text data, image-text data is easier to collect and ensures higher data quality. For vision-centric framework designs, we propose adapting the vision encoder to handle images of any resolution during the image understanding enhancement stage and tuning the encoder to efficiently embed video inputs.
@@ 2.1 Any-resolution Vision Tokenization
In MLLMs, visual inputs are extracted into vision tokens for multimodal understanding. The common practice liu2023visualit ; liu2023improved is to extract visual inputs with a pre-trained ViT-based vision encoder. The pre-trained vision encoder only receives images with fixed resolutions, which introduces information loss. To alleviate information loss, AnyRes techniques li2024llavaonevision ; internvl2.5 ; chen2023internvl are proposed to split images into patches with fixed resolutions. Although AnyRes techniques increase the number of vision tokens, it is still inflexible and neglects the position relationship within an image when extracting vision tokens. In VideoLLaMA3, we adopt the idea of Any-resolution Vision Tokenization (AVT) dehghani2023patch ; qwen2-vl to dynamically process images and videos of any resolution. Concretely, we adapt the pre-trained vision encoder (ViT-based architectures) to handle variable resolutions by employing a strategy to replace the absolute position embeddings in ViT with 2D-RoPE su2024roformer . With AVT, images and videos of different resolutions are better represented with more details included in vision tokens. To make the vision encoder compatible with AVT, we fine-tune the vision encoder and the projector in the stage of Vision Encoder Adaptation (i.e., stage #1 in Figure 2) using scene images, document data, and scene images with texts.
@@ 2.2 Differential Frame Pruner
For videos, inputs which usually have much more tokens than image inputs after tokenization, to reduce the computation demand for videos, we apply a per-frame 2$\times$2 spatial downsampling by bilinear interpolation to limit the context length within a certain range. Besides, considering that videos consist of frames with overlapping content, representing videos by stacking vision tokens from each frame leads to lengthy and redundant tokens. To further reduce the number of tokens of videos, we propose the Differential Frame Pruner (DiffFP) to prune the video tokens. Inspired by RLT (choudhury2024don, ), we compare the 1-norm distance between temporally consecutive patches within the pixel space. We consider temporally consecutive patches with smaller distances to be redundant, and the later patches can be pruned. Specifically, as shown in Figure 4, we first calculate the 1-norm distance between consecutive frames in the pixel space and then remove patches whose distances fall below a pre-defined threshold. Following RLT (choudhury2024don, ), we set the default threshold to 0.1.
@@ 3 Training
Inspired by previous explorations in MLLMs qwen2-vl ; li2024llavaonevision ; zhang2024video , we develop video understanding capabilities based on strong image understanding foundations. To enable the model with strong image and video understanding capabilities simultaneously, the training of VideoLLaMA3 has four stages: 1) Vision Encoder Adaptation, 2) Vision-Language Alignment, 3) Massive Multi-task Fine-tuning, and 4) Video-centric Fine-tuning. While the first three stages primarily focus on improving image understanding, the final stage is dedicated to video understanding. The details of the training stages are as follows:
1) Vision Encoder Adaptation Stage. In this stage, we fine-tune the vision encoder, which is initialized with the pre-trained SigLIP (Zhai2023SigmoidLF, ), on a large-scale image dataset. During this stage, the vision encoder is made trainable, while the language decoder remains frozen. This fine-tuning transforms the encoder into a dynamic-resolution processor, enhancing its ability to process images of varying resolutions. Meanwhile, the projector is trained to better align the features of the vision encoder with those of the LLM.
2)Vision-Language Alignment Stage. This stage primarily focuses on introducing multimodal knowledge into the model. During this phase, all parameters are made trainable, enabling both the LLM and the vision encoder to be fine-tuned for integrating multimodal knowledge.
3) Multi-task Fine-tuning Stage. In this stage, we perform instruction fine-tuning using a diverse set of multimodal question-answering data, which includes both image and video-based questions. This step is crucial for improving the model’s ability to follow natural language instructions and enhancing its multimodal understanding. Moreover, this stage lays the foundation for the model’s video understanding capabilities, enabling it to process and analyze temporal information. Also, in this stage, we introduce the video compressor to reduce the number of video tokens.
4) Video-centric Fine-tuning Stage. In this stage, we focus on enhancing the model’s video understanding capabilities. All parameters are unfreezed during this stage. The data used in this stage includes video-text data, image-only data and text-only data.
@@ 3.2.4 Video-centric Fine-tuning
In addition, we further expand the data scale and strengthen the model by synthesizing dense captions and QAs of specific aspects. Specifically, following the pipeline proposed in (zhang2024video, ), we first filter 68K dynamic videos from Panda-70M (chen2024panda, ) dataset by optical flow, and then employ Qwen2-VL-72B (qwen2-vl, ) to generate diverse dense captions and QAs for each video from the aspects of temporal understanding, spatial understanding, object description, and time-order understanding. Finally, 242K question-answer pairs are used for training.
@@ 4.1.1 Baselines
To comprehensively evaluate the image performance of VideoLLaMA3, we compare it against a diverse set of baselines. For the 2B version of the model, we select several strong methods, including SmolVLM (smolvlm2023, ), InternVL2.5-2B (internvl2.5, ), and Qwen2VL-2B (qwen2-vl, ). For the 7B model, there are more options available. We choose to compare against Molmo-7B-D (molmo2024, ), InternVL2.5-8B (internvl2.5, ), LLaVA-OneVision (li2024llavaonevision, ), NVILA (liu2024nvila, ), and Qwen2VL-8B (qwen2-vl, ).
@@ 4.1.2 Benchmarks
Finally, to evaluate VideoLLaMA3’s performance in general question answering, particularly in real-world and complex scenarios, we conduct assessments using several challenging benchmarks. The benchmarks include: 1) RealWorldQA (realworldqa, ), which focuses on answering questions based on realistic images drawn from everyday scenarios, 2) AI2D (kembhavi2016diagram, ), which evaluates the model’s ability to reason about diagrams and science images, 3) GQA (hudson2019gqa, ), which assesses general question answering with a focus on complex visual reasoning tasks, and 4) MME, which includes a wide variety of general knowledge questions that require a deep understanding of visual information.
@@ 4.2.1 Baselines
To comprehensively evaluate the video performance of VideoLLaMA3, we compare it with a diverse set of baseline models. Similar to image evaluation, there are few available models with a 2B parameter size in the community. We select several strong baselines, including Apollo-2B (zohar2024apollo, ), InternVL2.5-2B (internvl2.5, ), and Qwen2VL-2B (qwen2-vl, ). For the 7B model, we compare it with general models such as Qwen2VL-8B (qwen2-vl, ), InternVL2.5-8B (internvl2.5, ), and NVILA (liu2024nvila, ), as well as proprietary models like LLaVA-Video (zhang2024video, ), Apollo-7B (zohar2024apollo, ), and our previous generation model, VideoLLaMA2 (damonlpsg2024videollama2, ).
@@ 5 Related Work
Multimodal LLMs for Native Video Understanding. Early video MLLMs primarily relied on sparsely sampled frames and simple connectors, such as MLPs (maaz2023video, ; lin2023video, ; ataallah2024minigpt4, ), tokenizers (jin2024video, ), and Q-formers (zhang2023video, ; li2024mvbench, ), to link visual encoders with large language models. Subsequent models propose various methods to overcome token limitations and support long-form video understanding. For example, some (zhang2024longva, ) directly extend the context window of LLMs, while others (shen2024longvu, ; xu2024pllava, ; jin2024chat, ; zhang2024video, ; xu2024slowfast, ; weng2025longvlm, ; song2024moviellm, ; li2025llama, ; tan2024koala, ) introduce video token compression techniques that perform pooling across spatial, temporal, or both dimensions. While most approaches utilize image-based encoders (zohar2024apollo, ; song2024moviechat+, ; chen2024sharegpt4video, ; fei2024video, ; liu2024kangaroo, ; liu2024oryx, ), some incorporate video-specific encoders to better capture temporal dependencies (wang2024internvideo2, ; jung2024pegasus, ; Maaz2024VideoGPT+, ; cheng2024videollama, ).
More recent works (sun2024video, ; cheng2024videollama, ; ye2025cat, ; shu2023audio, ; su2023pandagpt, ; lyu2023macaw, ; panagopoulou2023x, ; han2024onellm, ; yu2024crema, ; tang2024avicuna, ) extend beyond visual inputs by incorporating audio, using separate encoders for each modality and integrating them through an LLM decoder. These models leverage joint instruction tuning on video-audio datasets (Yang2022AVQAAD, ; AlAmri2019AudioVS, ; ye2024cat, ) to capture interactions between visual and auditory information. Additionally, recent advances in streaming video understanding focus on real-time processing (zhang2024flash, ; qian2024streaming, ; Chen2024VideoLLMonlineOV, ; Zhang2024InternLMXComposer25OmniLiveAC, ; Liu2024StreamChatCW, ), employing techniques like adaptive memory and incremental processing for tasks such as live event detection and real-time captioning. Previous works shen2024longvu ; zhang2024video ; tan2024koala ; zohar2024apollo ; wang2024internvideo2 ; cheng2024videollama typically follow a training recipe that involves an alignment phase, followed by supervised fine-tuning, with instruction-tuning datasets (zhang2024video, ; Maaz2023VideoChatGPT, ; li2023mvbench, ; liu2024oryx, ; chen2024sharegpt4video, ) often being video dominant. However, we propose a vision-centric training paradigm to enhance video understanding capabilities by focusing on large-scale image understanding pre-training. This approach leverages high-quality image-text datasets to build robust vision encoders that are then adapted for video tasks.
@@
Furthermore, Qwen2-VL qwen2-vl adopts a unified framework for processing both images and videos, enhancing the model’s visual perception capabilities. Models such as Qwen2-VL, InternVL-2 chen2023internvl , and InternVL-2.5 internvl2.5 , which scale both model sizes (ranging from 1 billion to 78 billion parameters) and the volume of training data, have achieved highly competitive performance in both image and video understanding tasks. To address the challenges of processing longer video inputs, recent studies longllava ; internlmxcomposer25 have proposed solutions such as adapting model architectures by incorporating a hybrid design of Mamba and Transformer blocks or training with extensive long video datasets to support extended input and output sequences.

### v3

@@ Abstract
In this paper, we propose VideoLLaMA 3, a more advanced multimodal foundation model for image and video understanding. The core design philosophy of VideoLLaMA3 is vision-centric. The meaning of “vision-centric” is two-fold: the vision-centric training paradigm and vision-centric framework design. The key insight of our vision-centric training paradigm is that high-quality image-text data is crucial for both image and video understanding. Instead of preparing massive video-text datasets, we focus on constructing large-scale, high-quality image-text datasets. VideoLLaMA3 has four training stages: 1) Vision Encoder Adaptation, which enables the vision encoder to accept images of variable resolutions as input; 2) Vision-Language Alignment, which jointly tunes the vision encoder, projector, and LLM with large-scale image-text data covering multiple types (including scene images, documents, and charts) as well as text-only data. 3) Multi-task Fine-tuning, which incorporates image-text SFT data for downstream tasks and video-text data to establish a foundation for video understanding. 4) Video-centric Fine-tuning, which further improves the model’s capability in video understanding. As for the framework design, to better capture fine-grained details in images, the pretrained vision encoder is adapted to encode images of varying sizes into vision tokens with corresponding numbers, rather than a fixed number of tokens. For video inputs, we reduce the number of vision tokens according to their similarity so that the representation of videos will be more precise and compact. Benefiting from vision-centric designs, VideoLLaMA3 achieves compelling performances in both image and video understanding benchmarks.
@@ 1 Introduction
Figure 1: Performance Comparison of VideoLLaMA3 with the previous advanced image/video MLLM on various representative benchmarks. Specifically, VideoLLaMA3 not only demonstrates strong video understanding capabilities (VideoMME, PerceptionTest, MLVU) but also maintains excellent document comprehension abilities (DocVQA) and multimodal mathematical reasoning skills (MathVista). Note that LLaVA-OneVision is only used to evaluate image benchmarks, while LLaVA-Video is only used to evaluate video benchmarks.
@@
Existing MLLMs  li2023videochat ; lin2023video ; ataallah2024minigpt4 ; zhang2024longva ; zohar2024apollo ; song2024moviechat+ ; fei2024video ; liu2024kangaroo ; liu2024oryx ; wang2024internvideo2 ; jung2024pegasus ; Maaz2024VideoGPT+ ; cheng2024videollama ; sun2024video ; zhang2024video ; Maaz2023VideoChatGPT ; li2023mvbench ; chen2024sharegpt4video ; li2024llavaonevision ; mplug3 ; chen2023internvl ; longllava ; internlmxcomposer25 have made significant progress in multimodal understanding. Image-centric MLLMs wang2024qwen2 ; li2024llavaonevision ; internvl2.5 ; chen2023internvl ; molmo2024 ; liu2024nvila ; smolvlm2023 , leveraging high-quality image-text datasets li2024llava ; li2024llavaonevision ; tong2024cambrian ; mathew2021docvqadatasetvqadocument ; laurençon2024building ; masry2022chartqa ; huang-etal-2024-chart that are easier to collect and curate, have demonstrated strong performance in image understanding, such as visual question answering, OCR, and document understanding. Beyond static content such as images, video-centric MLLMs zhang2024video ; chen2024sharegpt4video ; cheng2024videollama ; damonlpsg2023videollama must tackle the added complexity of modeling the temporal dimension of videos, requiring models to handle dynamic content and capture dependencies across frames. This temporal complexity, combined with the need for large-scale video-text datasets that are often of lower quality and harder to annotate, makes video MLLMs more challenging. These challenges underscore the advantages of using image understanding as a foundation for video understanding. By extending the robust visual capabilities of image MLLMs, video models can focus on and better address the unique challenges of temporal and dynamic content modeling.
@@
Inheriting from VideoLLaMA DBLP:conf/emnlp/ZhangLB23 and VideoLLaMA2 damonlpsg2024videollama2 , VideoLLaMA3, a more advanced multimodal foundation model, is proposed for image and video understanding. We design VideoLLaMA3 in a vision-centric way. Specifically, we propose a vision-centric training paradigm and vision-centric framework designs. For the training paradigm, considering the intrinsic relationship between image and video modalities - where videos are essentially sequences of temporally correlated images, we prioritize the improvement of image understanding, which in turn enhances the performance of video understanding. Moreover, compared to video-text data, image-text data is easier to collect and ensures higher data quality. For vision-centric framework designs, we propose adapting the vision encoder to handle images of any resolution during the image understanding enhancement stage and tuning the encoder to efficiently embed video inputs.
@@ 2.1 Any-resolution Vision Tokenization
In MLLMs, visual inputs are extracted into vision tokens for multimodal understanding. The common practice liu2023visualit ; liu2023improved is to extract visual inputs with a pre-trained ViT-based vision encoder. The pre-trained vision encoder only receives images with fixed resolutions, which introduces information loss. To alleviate information loss, AnyRes techniques li2024llavaonevision ; internvl2.5 ; chen2023internvl are proposed to split images into patches with fixed resolutions. Although AnyRes techniques increase the number of vision tokens, it is still inflexible and neglects the position relationship within an image when extracting vision tokens. In VideoLLaMA3, we adopt the idea of Any-resolution Vision Tokenization (AVT) dehghani2023patch ; wang2024qwen2 to dynamically process images and videos of any resolution. Concretely, we adapt the pre-trained vision encoder (ViT-based architectures) to handle variable resolutions by employing a strategy to replace the absolute position embeddings in ViT with 2D-RoPE su2024roformer . With AVT, images and videos of different resolutions are better represented with more details included in vision tokens. To make the vision encoder compatible with AVT, we fine-tune the vision encoder and the projector in the stage of Vision Encoder Adaptation (i.e., stage #1 in Figure 2) using scene images, document data, and scene images with texts.
@@ 2.2 Differential Frame Pruner
For videos, inputs usually have much more tokens than image inputs after tokenization. To reduce the computation demand for videos, we apply a per-frame 2$\times$2 spatial downsampling by bilinear interpolation to limit the context length within a certain range. Besides, considering that videos consist of frames with overlapping content, representing videos by stacking vision tokens from each frame leads to lengthy and redundant tokens. To further reduce the number of tokens of videos, we propose the Differential Frame Pruner (DiffFP) to prune the video tokens. Inspired by RLT (choudhury2024don, ), we compare the 1-norm distance between temporally consecutive patches within the pixel space. We consider temporally consecutive patches with smaller distances to be redundant, and the later patches can be pruned. Specifically, as shown in Figure 4, we first calculate the 1-norm distance between consecutive frames in the pixel space and then remove patches whose distances fall below a pre-defined threshold. Following RLT (choudhury2024don, ), we set the default threshold to 0.1.
@@ 3 Training
Inspired by previous explorations in MLLMs wang2024qwen2 ; li2024llavaonevision ; zhang2024video , we develop video understanding capabilities based on strong image understanding foundations. To enable the model with strong image and video understanding capabilities simultaneously, the training of VideoLLaMA3 has four stages: 1) Vision Encoder Adaptation, 2) Vision-Language Alignment, 3) Multi-task Fine-tuning, and 4) Video-centric Fine-tuning. While the first three stages primarily focus on improving image understanding, the final stage is dedicated to video understanding. The details of the training stages are as follows:
1) Vision Encoder Adaptation. In this stage, we fine-tune the vision encoder, which is initialized with the pre-trained SigLIP (Zhai2023SigmoidLF, ), on a large-scale image dataset. During this stage, the vision encoder is made trainable, while the language decoder remains frozen. This fine-tuning transforms the encoder into a dynamic-resolution processor, enhancing its ability to process images of varying resolutions. Meanwhile, the projector is trained to better align the features of the vision encoder with those of the LLM.
2)Vision-Language Alignment. This stage primarily focuses on introducing multimodal knowledge into the model. During this phase, all parameters are made trainable, enabling both the LLM and the vision encoder to be fine-tuned for integrating multimodal knowledge.
3) Multi-task Fine-tuning. In this stage, we perform instruction fine-tuning using a diverse set of multimodal question-answering data, which includes both image and video-based questions. This step is crucial for improving the model’s ability to follow natural language instructions and enhancing its multimodal understanding. Moreover, this stage lays the foundation for the model’s video understanding capabilities, enabling it to process and analyze temporal information. Also, in this stage, we introduce the video compressor to reduce the number of video tokens.
4) Video-centric Fine-tuning. In this stage, we focus on enhancing the model’s video understanding capabilities. All parameters are unfreezed during this stage. The data used in this stage includes video-text data, image-only data and text-only data.
@@ 3.2.4 Video-centric Fine-tuning
In addition, we further expand the data scale and strengthen the model by synthesizing dense captions and QAs of specific aspects. Specifically, following the pipeline proposed in (zhang2024video, ), we first filter 68K dynamic videos from Panda-70M (chen2024panda, ) dataset by optical flow, and then employ Qwen2-VL-72B (wang2024qwen2, ) to generate diverse dense captions and QAs for each video from the aspects of temporal understanding, spatial understanding, object description, and time-order understanding. Finally, 242K question-answer pairs are used for training.
@@ 4.1.1 Baselines
To comprehensively evaluate the image performance of VideoLLaMA3, we compare it against a diverse set of baselines. For the 2B version of the model, we select several strong methods, including SmolVLM (smolvlm2023, ), InternVL2.5-2B (internvl2.5, ), and Qwen2VL-2B (wang2024qwen2, ). For the 7B model, there are more options available. We choose to compare against Molmo-7B-D (molmo2024, ), InternVL2.5-8B (internvl2.5, ), LLaVA-OneVision (li2024llavaonevision, ), NVILA (liu2024nvila, ), and Qwen2VL-8B (wang2024qwen2, ).
@@ 4.1.2 Benchmarks
Finally, to evaluate VideoLLaMA3’s performance in general question answering, particularly in real-world and complex scenarios, we conduct assessments using several challenging benchmarks. The benchmarks include: 1) RealWorldQA (realworldqa, ), which focuses on answering questions based on realistic images drawn from everyday scenarios, 2) AI2D (kembhavi2016diagram, ), which evaluates the model’s ability to reason about diagrams and science images, 3) GQA (hudson2019gqa, ), which assesses general question answering with a focus on complex visual reasoning tasks, and 4) MME (fu2023mme, ), which includes a wide variety of general knowledge questions that require a deep understanding of visual information.
@@ 4.2.1 Baselines
To comprehensively evaluate the video performance of VideoLLaMA3, we compare it with a diverse set of baseline models. Similar to image evaluation, there are few available models with a 2B parameter size in the community. We select several strong baselines, including Apollo-2B (zohar2024apollo, ), InternVL2.5-2B (internvl2.5, ), and Qwen2VL-2B (wang2024qwen2, ). For the 7B model, we compare it with generalist models such as Qwen2VL-7B (wang2024qwen2, ), InternVL2.5-8B (internvl2.5, ), and NVILA (liu2024nvila, ), as well as specialist models like LLaVA-Video (zhang2024video, ), Apollo-7B (zohar2024apollo, ), and our previous generation model, VideoLLaMA2 (damonlpsg2024videollama2, ).
@@ 5 Related Work
Multimodal LLMs for Native Video Understanding. Early video MLLMs primarily relied on sparsely sampled frames and simple connectors, such as MLPs (maaz2023video, ; lin2023video, ; ataallah2024minigpt4, ), discrete visual tokenizers (jin2024video, ), and Q-formers (zhang2023video, ; li2024mvbench, ), to link visual encoders with large language models. Subsequent models propose various methods to overcome token limitations and support long-form video understanding. For example,  (zhang2024longva, ) directly extends the context window of LLMs to achieve long video understanding, while others (shen2024longvu, ; xu2024pllava, ; jin2024chat, ; zhang2024video, ; xu2024slowfast, ; weng2025longvlm, ; song2024moviellm, ; li2025llama, ; tan2024koala, ) introduce video token compression techniques that perform pooling across spatial, temporal, or both dimensions. While most approaches utilize image-based encoders (zohar2024apollo, ; song2024moviechat+, ; chen2024sharegpt4video, ; fei2024video, ; wang2024tarsierrecipestrainingevaluating, ; hong2024cogvlm2, ; liu2024kangaroo, ; liu2024oryx, ; chen2024timemarker, ), some incorporate video-specific encoders to better capture temporal dependencies (wang2024internvideo2, ; jung2024pegasus, ; Maaz2024VideoGPT+, ; cheng2024videollama, ).
More recent works (sun2024video, ; cheng2024videollama, ; ye2025cat, ; shu2023audio, ; su2023pandagpt, ; lyu2023macaw, ; panagopoulou2023x, ; han2024onellm, ; yu2024crema, ; tang2024avicuna, ) extend beyond visual inputs by incorporating audio, using separate encoders for each modality and integrating them through an LLM decoder. These models leverage joint instruction tuning on video-audio datasets (Yang2022AVQAAD, ; AlAmri2019AudioVS, ; ye2024cat, ; leng2024curse, ) to capture interactions between visual and auditory information. Additionally, recent advances in streaming video understanding focus on real-time processing (zhang2024flash, ; qian2024streaming, ; Chen2024VideoLLMonlineOV, ; Zhang2024InternLMXComposer25OmniLiveAC, ; Liu2024StreamChatCW, ), employing techniques like adaptive memory and incremental processing for tasks such as live event detection and real-time captioning. Previous works shen2024longvu ; zhang2024video ; tan2024koala ; zohar2024apollo ; wang2024internvideo2 ; cheng2024videollama typically follow a training recipe that involves an alignment phase, followed by supervised fine-tuning, with instruction-tuning datasets (zhang2024video, ; Maaz2023VideoChatGPT, ; li2023mvbench, ; liu2024oryx, ; chen2024sharegpt4video, ) often being video dominant. However, we propose a vision-centric training paradigm to enhance video understanding capabilities by focusing on large-scale image understanding pre-training. This approach leverages high-quality image-text datasets to build robust vision encoders that are then adapted for video tasks.
@@
Furthermore, Qwen2-VL wang2024qwen2 adopts a unified framework for processing both images and videos, enhancing the model’s visual perception capabilities. Models such as Qwen2-VL, InternVL-2 chen2023internvl , and InternVL-2.5 internvl2.5 , which scale both model sizes (ranging from 1 billion to 78 billion parameters) and the volume of training data, have achieved highly competitive performance in both image and video understanding tasks. To address the challenges of processing longer video inputs, recent studies longllava ; internlmxcomposer25 have proposed solutions such as adapting model architectures by incorporating a hybrid design of Mamba and Transformer blocks or training with extensive long video datasets to support extended input and output sequences.